a particle executin shm covers a distance x in first second from its mean position and it covers distance y in another one second. find amplitude of the particle

Grade:12

1 Answers

AJIT AskiitiansExpert-IITD
13 years ago

Dear aabhas ,

the genreal equation of shm is     x(t) =  Asin(wt)

so just plug in the values given in the quesiton and solve ....x

x = A sinw ------(1)

y = Asin2w - Asinw

y =2Acosw *sinw - Asinw

y = (2cosW -1) *x

cosw =(y/x+1)/2 =  (x+y)/2x

so sinw  = (1-cos2w)1/2  = ( 1 -(x+y/2x)2)1/2  =  (3x2-y2-2xy)1/2 / 2x

putting in eqn 1

A = x/sinw  = 2x2/(3x2-y2-2xy)1/2   

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Win exciting gifts by answering the questions on Discussion Forum